# New versions of Frobenius and integral closure of ideals

### v2

@@ 4.3.7. The Vraciu Frobenius closure
By Theorem 4.5, a sufficient condition for $x^{\mathbf{v}}$ to be in ${}^{\mathfrak{a}^{t}}I^{F}$ is for $\mathbf{v}+tH\subset\mathcal{B}+sH+\sigma^{\vee}$ for some $s>t$. Letting $s=t+\epsilon$ for some small $\epsilon>0$, we have that this condition is satisfied for any $\mathbf{v}$ with $x^{\mathbf{v}}\in I$ or $v_{1}\geq\max\{4+2\epsilon,6-2t\}$. Thus, we have that if $t\geq\frac{1}{2}$, then ${}^{\mathfrak{a}^{t}}I^{F}\supseteq I+(x^{5}y^{2})$.
Now, if $t<\frac{1}{2}$, then $I\subset{}^{\mathfrak{a}^{t}}I^{F}\subset{}^{\mathfrak{a}^{t}}I^{*}=I$, so ${}^{\mathfrak{a}^{t}}I^{F}=I$. If $\frac{1}{2}\leq t<1$, then $I+(x^{5}y^{2})\subset{}^{\mathfrak{a}^{t}}I^{F}\subset{}^{\mathfrak{a}^{t}}I^{*}=I+(x^{5}y^{2})$, and so ${}^{\mathfrak{a}^{t}}I^{F}=I+(x^{5}y^{2})$. If $t\geq 1$, then $I+(x^{5}y^{2})\subset{}^{\mathfrak{a}^{t}}I^{F}\subset{}^{\mathfrak{a}^{t}}I^{*}=I+(x^{4}y,x^{4}y^{2})$.
To determine ${}^{\mathfrak{a}^{t}}I^{F}$ for $t\geq 1$, we use the definition to check whether the monomials $x^{4}y$ and $x^{4}y^{2}$ belong to ${}^{\mathfrak{a}^{t}}I^{F}$. For a power $q$ of $p$ we have that
$(x^{4}y)^{q}\mathfrak{a}^{\lceil tq\rceil}=(x^{4}y)^{q}(x^{2},x^{2}y^{3})^{\lceil tq\rceil}=(x^{4q+2\lceil tq\rceil}y^{q+3i}\mid 0\leq i\leq\lceil tq\rceil)$
and
$\mathfrak{a}^{\lceil tq\rceil}I^{[q]}=(x^{2},x^{2}y^{3})^{\lceil tq\rceil}(x^{4q},x^{2q}y^{3q})=(x^{4q+2\lceil tq\rceil}y^{3j},x^{2q+2\lceil tq\rceil}y^{3q+3j}\mid 0\leq j\leq\lceil tq\rceil).$
Now $(x^{4}y)^{q}\mathfrak{a}^{\lceil tq\rceil}\subset\mathfrak{a}^{\lceil tq\rceil}I^{[q]}$ if and only if for every $i$, $0\leq i\leq\lceil tq\rceil$, there exists $j$, $0\leq j\leq\lceil tq\rceil$, such that $x^{4q+2\lceil tq\rceil}y^{q+3i}\in(x^{4q+\lceil tq\rceil}y^{3j})$ or $x^{4q+2\lceil tq\rceil}y^{q+3i}\in(x^{2q+\lceil tq\rceil}y^{3q+3j})$. Now, in $R$, we have that $x^{a}y^{b}\in(x^{c}y^{d})$ if and only if $(a,b)-(c,d)\in S$, which occurs exactly when $0\leq 2(b-d)\leq 3(a-c)$. Therefore,
$x^{4q+2\lceil tq\rceil}y^{q+3i}\in(x^{4q+2\lceil tq\rceil}y^{3j})\iff 0\leq 2(q+3i-3j)\leq 3(4q+2\lceil tq\rceil-(4q+2\lceil tq\rceil))=0.$
Thus, $x^{4q+2\lceil tq\rceil}y^{q+3i}\in(x^{4q+2\lceil tq\rceil}y^{3j})$ if and only if $q+3i-3j=0$. This is impossible unless $p=3$, in which case $j=\frac{q+3i}{3}$. Since $j\leq\lceil tq\rceil$, this is satisfiable so long as $q+3i\leq 3\lceil tq\rceil$, that is, $i\leq\lceil tq\rceil-\frac{q}{3}$.
On the other hand,
$x^{4q+2\lceil tq\rceil}y^{q+3i}\in(x^{2q+2\lceil tq\rceil}y^{3q+3j})\iff 0\leq 2(q+3i-(3q+3j))\leq 3(4q+2\lceil tq\rceil-(2q+2\lceil tq\rceil)),$
which simplifies to
$0\leq 6i-6j-4q\leq 6q\implies i-\frac{5}{3}q\leq j\leq i-\frac{2}{3}q.$
There exists a $j$ satisfying this so long as $i-\frac{2}{3}q\geq 0$, since $i-\frac{5}{3}q\leq i\leq\lceil tq\rceil$.
Thus, we have that $x^{4q+2\lceil tq\rceil}y^{q+3i}\in(x^{4q+2\lceil tq\rceil}y^{3j})$ for some $j$ if and only if $p=3$ and $i\leq\lceil tq\rceil-\frac{q}{3}$, and we have that $x^{4q+2\lceil tq\rceil}y^{q+3i}\in(x^{2q+2\lceil tq\rceil}y^{3q+3j})$ for some $j$ if and only if $i\geq\frac{2}{3}q$. In order for at least one of these containments to hold for each $0\leq i\leq\lceil tq\rceil$, we must have that $p=3$ and $\frac{2}{3}q\leq\lceil tq\rceil-\frac{q}{3}$, i.e. $q\leq\lceil tq\rceil$. This holds for all $q\gg 0$ if and only if $t\geq 1$.
Working similarly to check $x^{4}y^{2}$, we have that
$(x^{4}y^{2})^{q}\mathfrak{a}^{\lceil tq\rceil}=(x^{4}y^{2})^{q}(x^{2},x^{2}y^{3})^{\lceil tq\rceil}=(x^{4q+2\lceil tq\rceil}y^{2q+3i}\mid 0\leq i\leq\lceil tq\rceil)$
and as before
$\mathfrak{a}^{\lceil tq\rceil}I^{[q]}=(x^{4q+2\lceil tq\rceil}y^{3j},x^{2q+2\lceil tq\rceil}y^{3q+3j}\mid 0\leq j\leq\lceil tq\rceil).$
Now $(x^{4}y^{2})^{q}\mathfrak{a}^{\lceil tq\rceil}\subset\mathfrak{a}^{\lceil tq\rceil}I^{[q]}$ if and only if for every $i$, $0\leq i\leq\lceil tq\rceil$, there exists $j$, $0\leq j\leq\lceil tq\rceil$, such that $x^{4q+2\lceil tq\rceil}y^{2q+3i}\in(x^{4q+\lceil tq\rceil}y^{3j})$ or $x^{4q+2\lceil tq\rceil}y^{q+3i}\in(x^{2q+\lceil tq\rceil}y^{3q+3j})$. We have that
$x^{4q+2\lceil tq\rceil}y^{2q+3i}\in(x^{4q+2\lceil tq\rceil}y^{3j})\iff 0\leq 2(2q+3i-3j)\leq 3(4q+2\lceil tq\rceil-(4q+2\lceil tq\rceil))=0.$
Thus, $x^{4q+2\lceil tq\rceil}y^{2q+2i}\in(x^{4q+2\lceil tq\rceil}y^{2j})$ if and only if $02q+3i-3j=$ 0, which implies $j=\frac{2q+3i}{3}$. This is again only possible if $p=3$, in which case, we can find such a $j$ so long as $2q+3i\leq 3\lceil tq\rceil$, that is, $i\leq\lceil tq\rceil-\frac{2}{3}q$.
On the other hand,
$x^{4q+2\lceil tq\rceil}y^{2q+3i}\in(x^{2q+2\lceil tq\rceil}y^{3q+3j})\iff 0\leq 2(2q+3i-(3q+3j))\leq 3(4q+2\lceil tq\rceil-(2q+2\lceil tq\rceil)),$
which simplifies to
$0\leq 6i-6j-2q\leq 6q\implies i-\frac{4}{3}q\leq j\leq i-\frac{1}{3}q.$
There exists a $j$ satisfying this so long as $i-\frac{1}{3}q\geq 0$, since $i-\frac{4}{3}q\leq i\leq\lceil tq\rceil$.
Thus, we have that $x^{4q+2\lceil tq\rceil}y^{2q+3i}\in(x^{4q+2\lceil tq\rceil}y^{3j})$ for some $j$ if and only if $p=3$ and $i\leq\lceil tq\rceil-\frac{2}{3}q$, and we have that $x^{4q+2\lceil tq\rceil}y^{2q+3i}\in(x^{2q+2\lceil tq\rceil}y^{3q+3j})$ for some $j$ if and only if $i\geq\frac{1}{3}q$. In order for at least one of these containments to hold for each $0\leq i\leq\lceil tq\rceil$, we must have that $p=3$ and $\frac{1}{3}q\leq\lceil tq\rceil-\frac{2}{3}q$, i.e. $q\leq\lceil tq\rceil$. This holds for all $q\gg 0$ if and only if $t\geq 1$.
@@
So, we have found that ${}^{\mathfrak{a}^{t}}I^{F}$ for $t\geq 1$ depends on $p$, and hence cannot be detected simply using the convex geometry of the affine semigroup. Finally,
${}^{\mathfrak{a}^{t}}I^{F}=\begin{cases}K_{1}=I&\text{ if }t<\frac{1}{2}\\ K_{2}=K_{1}+(x^{5}y^{2})=(x^{4},x^{5}y^{2},x^{2}y^{3})&\text{ if }t\geq\frac{1}{2},\text{ and }t<1\text{ or }p\neq 3\\ K_{3}=K_{2}+(x^{4}y,x^{4}y^{2})=(x^{4},x^{4}y,x^{4}y^{2},x^{2}y^{3})&\text{ if}t\geq 1\text{ and }p=3\end{cases}$
4.4. A 3-generated example
Let $R=k[x,y]$, $\mathfrak{a}=(xy^{3},x^{2}y,x^{3})$, and $I=(x^{4}y,xy^{4})$, so that $\mathbf{b}_{i}=(7-3i,3i-2)$ for $i=1,2$. We wish to compute the various $\mathfrak{a}^{t}$-closures of $I$. In this case we have that $\Delta=\{(\alpha,\beta,\gamma)^{T}\mid\alpha+\beta+\gamma=1,\alpha,\beta,\gamma\geq 0\}$.
For $\mathbf{u}=(\alpha,\beta,\gamma)^{T}\in\Delta$ and $\mathbf{v}=(v_{1},v_{2})^{T}$, we have that
$\mathbf{v}+tA\mathbf{u}\in\mathbf{\mathcal{missing}}{B}+\sigma^{\vee}$
$\iff\exists i\in\{1,2\}:\quad C^{T}(\mathbf{v}-\mathbf{b}_{i}+tA\mathbf{u})\geq\mathbf{0}$
$\iff\exists i\in\{1,2\}:\quad\left(\begin{array}[]{cc}0&1\\ 1&0\end{array}\right)\left(\left(\begin{array}[]{c}v_{1}\\ v_{2}\end{array}\right)-\left(\begin{array}[]{c}7-3i\\ 3i-2\end{array}\right)+t\left(\begin{array}[]{ccc}1&2&3\\ 3&1&0\end{array}\right)\left(\begin{array}[]{c}\alpha\\ \beta\\ \gamma\end{array}\right)\right)\geq\mathbf{0}$
$\iff\exists i\in\{1,2\}:\quad\left(\begin{array}[]{c}v_{2}-3i+2\\ v_{1}-7+3i\end{array}\right)+t\left(\begin{array}[]{c}3\alpha+\beta\\ \alpha+2\beta+3\gamma\end{array}\right)\geq\mathbf{0}$
$\iff\exists i\in\{1,2\}:\quad t\left(\begin{array}[]{c}3\alpha+\beta\\ \alpha+2\beta+3\gamma\end{array}\right)\geq\left(\begin{array}[]{c}3i-2-v_{2}\\ 7-3i-v_{1}\end{array}\right)$
$\iff\exists i\in\{1,2\}:\quad t\geq\max\left(\frac{3i-2-v_{2}}{3\alpha+\beta},\frac{7-3i-v_{1}}{\alpha+2\beta+3\gamma}\right)$
$\iff t\geq\min\left(\max\left(\frac{1-v_{2}}{3\alpha+\beta},\frac{4-v_{1}}{\alpha+2\beta+3\gamma}\right),\max\left(\frac{4-v_{2}}{3\alpha+\beta},\frac{1-v_{1}}{\alpha+2\beta+3\gamma}\right)\right),$
where if $03\alpha+\beta=$ 0, we interpret $\frac{3i-2-v_{2}}{3\alpha+\beta}$ as $\infty$ if $03i-2-v_{2}>$ 0, $\infty$- ∞ if $03i-2-v_{2}<$ 0, and 0 if $03i-2-v_{2}=$ 0. Thus, $\mathbf{v}+tH\subset\mathcal{B}+\sigma^{\vee}$ if and only if for all $(\alpha,\beta,\gamma)^{T}\in\Delta$, the above inequality holds, that is,
$t\geq\sup_{(\alpha,\beta\gamma)^{T}\in\Delta}\left\{\min\left(\max\left(\frac{1-v_{2}}{3\alpha+\beta},\frac{4-v_{1}}{\alpha+2\beta+3\gamma}\right),\max\left(\frac{4-v_{2}}{3\alpha+\beta},\frac{1-v_{1}}{\alpha+2\beta+3\gamma}\right)\right)\right\}.$
By way of example, we find the smallest $t$ such that $x^{3}y^{2}\in I^{F\mathfrak{a}^{t}}$. This gives us $v_{1}=3$, $v_{2}=2$, and so we wish to compute
$\sup_{(\alpha,\beta\gamma)^{T}\in\Delta}\left\{\min\left(\max\left(\frac{1-v_{2}}{3\alpha+\beta},\frac{4-v_{1}}{\alpha+2\beta+3\gamma}\right),\max\left(\frac{4-v_{2}}{3\alpha+\beta},\frac{1-v_{1}}{\alpha+2\beta+3\gamma}\right)\right)\right\}$
$=$ $\sup_{(\alpha,\beta\gamma)^{T}\in\Delta}\left\{\min\left(\max\left(\frac{-1}{3\alpha+\beta},\frac{1}{\alpha+2\beta+3\gamma}\right),\max\left(\frac{2}{3\alpha+\beta},\frac{-2}{\alpha+2\beta+3\gamma}\right)\right)\right\}$
$=$ $\sup_{(\alpha,\beta\gamma)^{T}\in\Delta}\left\{\min\left(\frac{1}{\alpha+2\beta+3\gamma},\frac{2}{3\alpha+\beta}\right)\right\}.$
Now $\frac{2}{3\alpha+\beta}\leq\frac{1}{\alpha+2\beta+3\gamma}$ if and only if $\alpha-3\beta-6\gamma\geq 0$, and so the supremum above is equal to
$\sup\left\{\sup_{\alpha-3\beta-6\gamma\geq 0}\left\{\frac{2}{3\alpha+\beta}\right\},\sup_{\alpha-3\beta-6\gamma\leq 0}\left\{\frac{1}{\alpha+2\beta+3\gamma}\right\}\right\}=\sup\left\{\frac{4}{5},\frac{4}{5}\right\}=\frac{4}{5}.$
Therefore $x^{3}y^{2}\in I^{F\mathfrak{a}^{t}}$ so long as $t\geq\frac{4}{5}$.

### v3

@@ A.6. The Vraciu Frobenius closure
By Corollary 6.6, a sufficient condition for $x^{\mathbf{v}}$ to be in $^{\mathfrak{a}^{t}}I^{F}$ is for $\mathbf{v}+tH\subset\mathcal{B}+sH+\sigma^{\vee}$ for some $s>t$. Letting $s=t+\epsilon$ for some small $\epsilon>0$ in (A.11), we have that this condition is satisfied for any $\mathbf{v}$ with $x^{\mathbf{v}}\in I$ or $v_{1}\geq\max\{4+2\epsilon,6-2t\}$. Thus, we have that if $t\geq 1/2$, then $^{\mathfrak{a}^{t}}I^{F}\supseteq I+(x^{5}y^{2})$.
Now, if $t<1/2$, then $I\subset{}^{\mathfrak{a}^{t}}I^{F}\subset{}^{\mathfrak{a}^{t}}I^{*}=I$, so $^{\mathfrak{a}^{t}}I^{F}=I$. If $1/2\leq t<1$, then $I+(x^{5}y^{2})\subset{}^{\mathfrak{a}^{t}}I^{F}\subset{}^{\mathfrak{a}^{t}}I^{*}=I+(x^{5}y^{2})$, and so $^{\mathfrak{a}^{t}}I^{F}=I+(x^{5}y^{2})$. If $t\geq 1$, then $I+(x^{5}y^{2})\subset{}^{\mathfrak{a}^{t}}I^{F}\subset{}^{\mathfrak{a}^{t}}I^{*}=I+(x^{4}y,x^{4}y^{2})$.
To determine $^{\mathfrak{a}^{t}}I^{F}$ for $t\geq 1$, we use the definition to check whether the monomial $x^{4}y^{r}$ belongs to $^{\mathfrak{a}^{t}}I^{F}$ for $r\in\{1,2\}$.
If $x^{4}y^{r}\in{}^{\mathfrak{a}^{t}}I^{F}$, then since $x^{2}\in\mathfrak{a}$, there exists $e$ such that
(A.12) $x^{4p^{e}}y^{rp^{e}}x^{2\lceil tp^{e}\rceil}\in(x^{4}y)^{p^{e}}\mathfrak{a}^{\lceil tp^{e}\rceil}\subset(x^{4},x^{2}y^{3})^{[p^{e}]}(x^{2},x^{2}y^{3})^{\lceil tp^{e}\rceil}.$
Since comparing degrees of $y$ shows that $x^{4p^{e}}y^{rp^{e}}x^{2\lceil tp^{e}\rceil}\notin(x^{2}y^{3})^{[p^{e}]}$, if (A.12) holds then there exists $0\leq j\leq\lceil tp^{e}\rceil$ such that
$x^{4p^{e}+2\lceil tp^{e}\rceil}y^{rp^{e}}=x^{4p^{e}}y^{rp^{e}}x^{2\lceil tp^{e}\rceil}\in x^{4p^{e}}(x^{2\lceil tp^{e}\rceil}y^{3j})=(x^{4p^{e}+2\lceil tp^{e}\rceil}y^{3j}).$
Therefore $y^{rp^{e}-3j}\in R$. Since the only power of $y$ in $R$ is $1$, $rp^{e}=3j$, and so $p=3$.
On the other hand, suppose $p=3$ and let $e\geq 1$. The ideal $(x^{4}y^{r})^{3^{e}}\mathfrak{a}^{\lceil 3^{e}t\rceil}$ is generated by the monomials $(x^{4}y^{r})^{3^{e}}(x^{\lceil 3^{e}t\rceil}y^{3i})=x^{4\cdot 3^{e}+\lceil 3^{e}t\rceil}y^{3^{e}r+3i}$ for $0\leq i\leq\lceil 3^{e}t\rceil$.
If $i\leq 3^{e}-3^{e-1}r$, then let $j=3^{e-1}r+i\leq 3^{e}\leq\lceil 3^{e}t\rceil$, and note that
$x^{4\cdot 3^{e}+2\lceil 3^{e}t\rceil}y^{3^{e}r+3i}=(x^{4})^{3^{e}}x^{2\lceil 3^{e}t\rceil}y^{3^{e}r+3(j-3^{e-1}r)}=(x^{4})^{3^{e}}x^{2\lceil 3^{e}t\rceil}y^{3j}\in(x^{4},x^{2}y^{3})^{[3^{e}]}(x^{2},x^{2}y^{3})^{\lceil 3^{e}t\rceil}.$
If $i\geq 3^{e}-3^{e-1}r$, then let $j=i-3^{e}+3^{e-1}r\geq 0$, and note that
$x^{4\cdot 3^{e}+2\lceil 3^{e}t\rceil}y^{3^{e}r+3i}=(x^{2}y^{3})^{3^{e}}x^{2\cdot 3^{e}+2\lceil 3^{e}t\rceil}y^{3^{e}r+3i-3^{e+1}}=(x^{2}y^{3})^{3^{e}}x^{2\cdot 3^{e}+2\lceil 3^{e}t\rceil}y^{3^{e}r+3(j+3^{e}-3^{e-1}r)-3^{e+1}}$
$(x^{2}y^{3})^{3^{e}}x^{2\cdot 3^{e}}x^{2\lceil 3^{e}t\rceil}y^{3j}\in(x^{4},x^{2}y^{3})^{[3^{e}]}(x^{2},x^{2}y^{3})^{\lceil 3^{e}t\rceil}.$
Therefore $x^{4}y^{r}\in{}^{\mathfrak{a}^{t}}I^{F}$ in this case.
@@
So, we have found that $^{\mathfrak{a}^{t}}I^{F}$ for $t\geq 1$ depends on $p$. Namely, $^{\mathfrak{a}^{t}}I^{F}$ contains $x^{4}y$ and $x^{4}y^{2}$ if and only if $t\geq 1$ and $p=3$. Hence membership in $^{\mathfrak{a}^{t}}I^{F}$ cannot be detected simply using the convex geometry of the affine semigroup. This shows (6.5) in 6.10.